# Competition Dynamics Shape Algorithmic Phases of In-Context Learning

### v1

@@ 4.1 Linear Interpolation of Algorithms (LIA)
Here, $p_{\text{model}}$ and $p_{a}$ respectively denote the next-token predictions of the model and individual algorithms from set $\mathcal{A}$, given the sequence $\mathbf{x}_{1:t}$ as input; meanwhile, $w_{a}$ denotes the weight associated with algorithm $a\in\mathcal{A}$ in the interpolation. We optimize the interpolation weights by minimizing Eq. 7 over multiple ID sequences $\mathbf{x}_{1:t}$, i.e., sequences sampled from Markov chains that constitute $\mathcal{T}_{\text{train}}$ (see App. A.3.4 for further details). Fits are almost perfect for all settings (see App. H, Fig. 38).
@@ 4.3 Model Analysis Using Algorithmic Phase diagrams
  • Fig. 8 (c) shows the effect of expanding the token space to allow each token to represent the last state and transition (See App. A.2 for details). This experiment is motivated to understand the effect of a tokenization on downstream abilities. We find tokenization allows the model to be able to count transitions without a formation of a complex attention head, which we suspect is one of the reasons Bi-Ret is slower to learn than Uni-Ret (see App. I for a discussion and App. E.3 for attention head visualizations). As expected, we find the Uni-Ret phase disappears in this case, since Bi-Ret is both superior on the training set and (now) the simpler solution.
@@ 5 Conclusion
CFP and ESL independently defined the setup, ran preliminary investigations, and converged on a bulk of the experiments, evaluation protocols, and phenomenology. CFP ran the final experiments in Section 3 in collaboration with ESL to identify the algorithmic phases and uncover transience of ICL. This formed the basis of the paper. ESL proposed the LIA protocol in Section 4, which CFP and ESL used to develop explanations of ICL’s phenomenology; CFP led the experiments. ESL performed a first version of the mechanistic interpretability analysis, which IP significantly expanded. IP also proposed the memorization analysis and the analysis of attention maps at phase boundaries. ESL defined the project narrative and wrote the paper, with inputs from all authors. CFP, ESL, and HT designed the conceptual diagrams. HT supervised the project.
@@ C.1 List of phenomena we reproduce
    Formation of induction heads and variants: Induction heads are a specialized attention head that help infer next-token predictions in a context-conditioned manner Elhage et al. (2021); Olsson et al. (2022); Reddy (2023); Edelman et al. (2024); Akyürek et al. (2024). Often, there is a sudden loss drop that correlates with induction head formation, and we find consistently find this drop occurs in our results with training time: this drop, in fact, is the cause of sudden transition from unigram to bigram dependence in the model. We reproduce this phenomena as seen in Fig. 1 (b), Fig. 3 (a,c) and Fig. 27 (b). Interestingly, we find that data-diversity has almost no effect to the formation of this circuit. In App. J, we discuss that this transition is likely an “Optimization Limited Emergence” as opposed to the Bi-Inf vs. Bi-Ret transition which is mostly data limited.
@@ Appendix G Generalization to chains from a different prior
Appendix H Further validation of LIA
H.1 Does LIA fit the model well?
@@
H.2 Robustness of LIA to arbitrary algorithms
@@
Appendix I Importance of sequence-space structure to study transformers
@@
Appendix J Discussion and Future Directions
@@
Appendix K Code Availability
All code used to run experiment and analysis will be released after the peer review process.

### v3

@@ 4.1 Linear Interpolation of Algorithms (LIA)
Here, $p_{\text{model}}$ and $p_{a}$ respectively denote the next-token predictions of the model and individual algorithms from set $\mathcal{A}$, given the sequence $\mathbf{x}_{1:t}$ as input; meanwhile, $w_{a}$ denotes the weight associated with algorithm $a\in\mathcal{A}$ in the interpolation. We optimize the interpolation weights by minimizing Eq. 7 over multiple ID sequences $\mathbf{x}_{1:t}$, i.e., sequences sampled from Markov chains that constitute $\mathcal{T}_{\text{train}}$ (see App. A.3.4 for further details). Fits are almost perfect for all settings (see App. I, Fig. 39).
@@ 4.3 Model Analysis Using Algorithmic Phase diagrams
  • Fig. 8 (c) shows the effect of expanding the token space to allow each token to represent the last state and transition (See App. A.2 for details). This experiment is motivated to understand the effect of a tokenization on downstream abilities. We find tokenization allows the model to be able to count transitions without a formation of a complex attention head, which we suspect is one of the reasons Bi-Ret is slower to learn than Uni-Ret (see App. J for a discussion and App. E.3 for attention head visualizations). As expected, we find the Uni-Ret phase disappears in this case, since Bi-Ret is both superior on the training set and (now) the simpler solution.
@@ 5 Conclusion
CFP and ESL independently defined the setup, ran preliminary investigations, and converged on a bulk of the experiments, evaluation protocols, and phenomenology; CFP ran extensive experiments to verify the latter. CFP ran the final experiments in Section 3 in collaboration with ESL to identify the algorithmic phases and uncover transience of ICL. This formed the basis of the paper. ESL and CFP proposed the LIA protocol in Section 4, which they used to develop explanations for ICL’s phenomenology; CFP led the experiments. ESL and CFP performed a first version of the mechanistic interpretability analysis, which IP significantly expanded. IP also proposed the memorization analysis and the analysis of attention maps at phase boundaries. ESL defined the project narrative and wrote the paper, with inputs from all authors. CFP, ESL, and HT designed the conceptual diagrams. HT supervised the project.
@@ C.1 List of phenomena we reproduce
    Formation of induction heads and variants: Induction heads are a specialized attention head that help infer next-token predictions in a context-conditioned manner Elhage et al. (2021); Olsson et al. (2022); Reddy (2023); Edelman et al. (2024); Akyürek et al. (2024). Often, there is a sudden loss drop that correlates with induction head formation, and we find consistently find this drop occurs in our results with training time: this drop, in fact, is the cause of sudden transition from unigram to bigram dependence in the model. We reproduce this phenomena as seen in Fig. 1 (b), Fig. 3 (a,c) and Fig. 27 (b). Interestingly, we find that data-diversity has almost no effect to the formation of this circuit. In App. K, we discuss that this transition is likely an “Optimization Limited Emergence” as opposed to the Bi-Inf vs. Bi-Ret transition which is mostly data limited.
@@ Appendix G Generalization to chains from a different prior
Appendix H In-Context Learning of Markov Chains in Real Large Language Models
Real language models trained on complex patterns including human languages, code and structured documents are known to acquire general in context learning abilities (Brown et al., 2020). Thus, we expect these models to deliver non-trivial performance at learning to continue a Markov sequence given in-context even though they haven’t been explicitly trained on this task. Furthermore, Wei et al. (2023) showed that larger language models “do in-context learning differently”, rougly speaking, in a more context-dependent way. To this end (and also motivated by curiosity) we evaluate the ICL ability of Llama-3.1 (Touvron et al., 2023; Dubey et al., 2024) across model scales: 8B, 70B and 405B. Since LLMs are known to be prompt sensitive (Sclar et al., 2024), we test three different prompts which differ by how precisely the Markovian nature of the sequence is described. We use NNsight and NDIF (Fiotto-Kaufman et al., 2024) to run these experiments.
Fig. 38 illustrates the results of inputting a sequence from a Markov chain as tokens of digits from 00 to $9$. More precisely, we drew a transition matrix $T^{*}$ from the DGP in App. A.1, and generated 300 sequences to collect next token probabilities to construct $\hat{T}$ as in App. A.3. We used a separator token, “,”, and we constrained the outputs to the digit tokens.
In Fig. 38 (left) we used a simple prompt without mentioning “Markov process”. Interestingly, we found a performance in between Uni-Inf and Bi-Inf, demonstrating Llama is able to model the sequence better than purely predicting a random probability from the stationary distribution (histogram), without even being instructed that the sequence is Markovian. Furthermore, bigger models (405B $>$ 70B $>$ 8B) are able to model the Markov sequence better (lower KL).
Fig. 38 (center, right) respectively used a prompt which mentions the process is Markovian and a prompt which additionally and explicitly states that next state probabilities only depends on the last state. We find a significant KL divergence improvement on the $70B$ model when adding these specifications, revealing that LLMs can, to some extent parse a natural language instruction and reflect it into its sequence modeling process.
Figure 38: In-context learning Markov chains on Llama-3.1. We input sequences generated from the DGP in App. A.1 into LLama-3.1 8B, 70B and 405B and measure the KL divergence on the next state probability between the model predictions and the ground truth transition matrix. Each state is tokenized into a single token, from “0” to “9”. The tokens are separated by the “,” token to avoid the tokenizer merging states together. The output space was constrained to the state tokens. (left) Prompt 1, not explicitly stating the Markovian nature of the sequence. (center) Prompt 2, stating the Markovian nature of the sequence. (right) Prompt 3, additionally stating that the next state probability will only depend on the last state.
An interesting future direction would be to reverse engineer what algorithmic solutions are implemented by these models to generate the next state, e.g. by decomposing the logit outputs as in Sec. 4.1.
Appendix I Further validation of LIA
I.1 Does LIA fit the model well?
@@
I.2 Robustness of LIA to arbitrary algorithms
@@
Appendix J Importance of sequence-space structure to study transformers
@@
Appendix K Discussion and Future Directions
@@
Appendix L Code Availability
All code used to run the experiments and analysis are available at